1. The most popular method of pricing is _____ pricing.

a. breakeven

b. incremental-cost

c. cost-plus

d. full-cost

2. Prestige pricing objectives emphasize:

a. cost minimization.

b. quality and exclusivity.

c. revenue maximization.

d. sales maximization.

3. Global Diamond manufactures jewelry. This firm is planning to introduce a new necklace and is trying to determine how many units it must sell in order to break even. Fixed costs are $100,000 and variable costs for each unit will be $20. At the price of $45 each, the number of units that must be sold in order to break even is:

a. 5,000.

b. 7,500.

c. 2,500.

d. 4,000.
Answer #1

Q1 : Cost-plus pricing

Cost plus pricing is the most popular method of pricing as it is the easiest one to implement. Cost plus pricing means having a markup on top of all the costs. The firm as a producer of goods is aware of all the costs which will be incurred in the process of production. It also knows what profit it wants to make from the products. Hence both the components of the cost-plus pricing are easily known without any complex calculation and hence this is the most popular method of pricing.

Q2 : Quality and exclusivity

In prestige pricing, the seller tries to influence the buyer by creating an impression that owning a product is a matter of prestige. The product is portrayed to be of sublime quality and one which only a few can afford or buy. Example of such products include high-end cars which are very expensive and each and every feature is portrayed to be of the highest quality.

Q3 : 4000

Breakeven quantity is calculated as the ratio fo Fixed Price and Contribution Margin, where Contribution Margin is the difference between selling price and variable cost. In the given case the fixed price is 100000, Seeling Price is $45 and variable cost is $20. Hence breakeven quantity = 100000/(45-20) = 100000/25 = 4000 units